OpenStudy (anonymous):

im asking these next few questions so i can kno that i was right myself so please write the answer and the way u sloved the problem Simplify: (3x+2y)2=

OpenStudy (anonymous):

(3x+2y)^2

OpenStudy (anonymous):

6x+4y

OpenStudy (anonymous):

it would be better if you posted your answer and we could tell you that you are correct

OpenStudy (anonymous):

ok hold on ima write it out

OpenStudy (anonymous):

or (3x+2y)^2

OpenStudy (anonymous):

(3x+2y)(3x+2y)=6x^2+4y^2

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

9x^2+12xy+4y^2

OpenStudy (anonymous):

u mutiply it ??

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

FOIL 3x(3x)+3x(2y)+2y(3x)+2y(2y) and simplify and combine like terms

OpenStudy (anonymous):

Or use 3x(3x+2y)+2y(3x+2y)

OpenStudy (anonymous):

Now multiply each and every terms

OpenStudy (anonymous):

o ok so basicaly i mutipliy evrything by the second power :)

OpenStudy (anonymous):

so i got 9x^2+4y^2

OpenStudy (anonymous):

no, you are multiplying each term 3x(3x)+3x(2y)+2y(3x)+2y(2y) 9x^2 +6xy +6xy +4Y^2 9x^2+12xy +4y^2

OpenStudy (anonymous):

use this formula (a+b)^2 = a^2+2ab+b^2

OpenStudy (anonymous):

huh im confused i thought u had to used the 2nd power for each part of the equtaion

OpenStudy (anonymous):

3^2 means 3 * 3 so ( 3x+2y ) ^2 means ( 3x+2y ) ( 3x+2y )